A sum of Rs. 24 is to be divided between Ayesha and Ali such that Ayesha will receive twice as much as Ali. How much money will Ali receive?
Answer by Theo(13342)

let x equal the amount that ali gets.
then 2x equals the amount that ayesha gets.

x + 2x = 24.

combine like terms to get 3x = 24.

solve for x to get x = 24 / 3 = 8.

ali gets 8.
ayesha gets 16.

ayesha gets twice as much as ali.

16 + 8 = 24.

your solution is that ali gets rs. 8.